What current is required in the windings of a long solenoid that has 1350 turns uni- formly distributed over a length of 0.128 m in order to produce a magnetic field of magnitude 0.000341 T at the center of the solenoid? The permeability of free space is 4π×10−7T·m/A. Answer in units of mA.
Answer #1

N = Number of turns = 1350

L = 0.128 m

n = number of turns per unit length = N/L = 1350 / 0.128 = 10546.875

B = magnetic field = 0.000341 T

i = current
using the formula

B = uo ni

0.000341 = (12.56 x 10-7) (10546.875 ) i

i = 0.02575 A = 25.75 mA